OpenStudy (anonymous):

does anyone have any help integrating this. (y^2)/(36-(y^2))^(3/2)

OpenStudy (anonymous):

trig sub

OpenStudy (nenadmatematika):

try with y=6sint....you will get trigonometric integral...

OpenStudy (anonymous):

I dont know what to do....@Outkast3r09 and @nenadmetamatika

OpenStudy (anonymous):

y=6sin(t)=asin(t) dy=6cos(t)dy=acos(t)dy \[\sqrt{36-y^2}=\sqrt{36-36sin^2t}=\sqrt{36(1-sin^2(t)} = 6 \sqrt{1-sin^2t}=\[6\sqrt{cos^2t}\]

OpenStudy (anonymous):

\[=6\sqrt{cos^2t}=6cos(t)\]

OpenStudy (anonymous):

replace these all in your equation and you'll have integral that's much easier

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

should get \[\int{}{}\frac{36sin^2t}{(6cos(t)^{3/2}}*6cos(t)dt\]

OpenStudy (anonymous):

thats not one of the answers i have. there were four choices given.

OpenStudy (anonymous):

what are the answers

OpenStudy (anonymous):

not to know you can simplify it

OpenStudy (anonymous):

a. square root (36-y^2)-sin^-1 (y/6) +c b. 6y/sqrt(36-y^2)-sin^-1 (y) +c c. y/sqrt(36-y^2)+c d. y/sqrt (36-y^2)-sin^-1 (y/6) +c

OpenStudy (anonymous):

so ive used this website but i just dont understand it at all. http://www.wolframalpha.com/input/?i=y%5E2%2F(36-y%5E2)%5E(3%2F2)%20

OpenStudy (anonymous):

get rid of the 3/2 it should just be this

OpenStudy (anonymous):

why do you get rid of it?

OpenStudy (anonymous):

\[\int{}{}\frac{36sin^2t}{6cos(t)^3}*6cos(t)dt \]

OpenStudy (anonymous):

so the answer is the last one?

OpenStudy (anonymous):

yes you end up after simplifying \[\int{}{}\frac{36sin^2t}{36cos^2t}dt\]

OpenStudy (anonymous):

which is the same as \[36\int{}{}tan^2tdt\]

OpenStudy (anonymous):

and you have to draw a triangle to figure out the tangent?

OpenStudy (anonymous):

which means tangent is equal to y/(squareroot (36-y^2))

OpenStudy (anonymous):

and i meant 36/36 which is one \[\int{}{}tan^2tdt\]

OpenStudy (anonymous):

do you know how to solve this equation?

OpenStudy (anonymous):

the equation would be (sec^2 t)-1 then integrate?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

making it tan t-t +c

OpenStudy (anonymous):

yep then take you trig sub y=6sin(t) and solve for t

OpenStudy (anonymous):

plug in y/(sqrt 36-y^2) - sin^(-1) (y/6) +c

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

sweet thanks!!

OpenStudy (anonymous):